The order for Cephalexin 500 mg PO every 12 hours is written by the MD. The...

  1. The order for Cephalexin 500 mg PO every 12 hours is written by the MD. The pharmacy sends a bottle with the label reading Cephalexin 250 mg/10mL. How many tsp will need to be given with this AM dose?
Answer #1

ANSWER.

* Calculation of tsp of drug required.

          Formula = Desired dose/Dose in hand * Quantity.

     Here

          Desired dose(D) = 500 mg.

          Dose in hand (H) = 250 mg.

               Quantity(Q) = 10 ml

As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 500/250 * 10

                                                     = 2 * 10

                                                     = 20 ml.

                               1 tsp = 5 ml.

        So

          tsp(Teaspoon) of drug required = 20/5

                                                                   = 4 tsp.